find the linearization L(X) at x=a
A.f(X)=x2-2x+3 at a=2?

Grade:11

1 Answers

D jaya
8 years ago
 f(x)=x2-2x+3
f(a)=a2-2a+3
f(2)=2(2)-2(2)+3
f(2)=4-4+3
f(2)=3           (or)
f’(a)=2-2+3
f’(20=0+3
f’(x)=3
l(x)=3+3(x-2)
     =3+3x-6
     =3x-3
      x=1
theirfore f(x)=x2-2x+3=1.